# DuetGraph: Coarse-to-Fine Knowledge Graph Reasoning with Dual-Pathway Global-Local Fusion

### v1

@@ Abstract
Knowledge graphs (KGs) are vital for enabling knowledge reasoning across various domains. Recent KG reasoning methods that integrate both global and local information have achieved promising results. However, existing methods often suffer from score over-smoothing, which blurs the distinction between correct and incorrect answers and hinders reasoning effectiveness. To address this, we propose DuetGraph, a coarse-to-fine KG reasoning mechanism with dual-pathway global-local fusion. DuetGraph tackles over-smoothing by segregating—rather than stacking—the processing of local (via message passing) and global (via attention) information into two distinct pathways, preventing mutual interference and preserving representational discrimination. In addition, DuetGraph introduces a coarse-to-fine optimization, which partitions entities into high- and low-score subsets. This strategy narrows the candidate space and sharpens the score gap between the two subsets, which alleviates over-smoothing and enhances inference quality. Extensive experiments on various datasets demonstrate that DuetGraph achieves state-of-the-art (SOTA) performance, with up to an 8.7% improvement in reasoning quality and a 1.8$\times$ acceleration in training efficiency.
@@ 1 Introduction
Knowledge graphs (KGs) are structured representations of real-world entities and their relationships, widely applied in domains such as information retrieval [1, 2], recommendation systems [3, 4], materials science [5, 6], and biomedical research [7, 8]. However, existing KGs are often incomplete, missing certain factual information [9, 10], which limits their effectiveness in downstream applications. As a result, inferring and completing missing entity information through KG reasoning is essential.
@@
However, such a single stage stacking approach tends to result in the problem of score over-smoothing, where incorrect answers receive scores similar to correct ones (Figure $1$The $x$-axis denotes normalized score gap ($\frac{|\text{Score}_{\text{correct}}-\text{Score}_{\text{incorrect}}|}{std(\text{Score})}$) and $y$-axis indicates entity count per interval.), making them hard to distinguish. Accordingly, we summarize the problem into two core challenges. Challenge 1: Existing studies [21, 22] have shown that stacking either message-passing or attention layers individually deepens information propagation and aggravates the over-smoothing problem. When message-passing and attention layers are stacked together, these effects accumulate. Challenge 2: The discriminative capacity of single stage models is typically limited [23], as they generate the answer directly based on a one-shot reasoning. This deficiency in discrimination further exacerbates the over-smoothing phenomenon [24].
@@
Our contributions can be summarized as follows. 1) We propose DuetGraph, a novel KG reasoning framework to alleviate score over-smoothing in KG reasoning. Specifically, DuetGraph: a) utilizes a dual-pathway fusion of local and global information instead of a single-pathway method, and b) adopts a coarse-to-fine design rather than one stage design. 2) We theoretically demonstrate that our proposed dual-pathway reasoning model and coarse-to-fine optimization can both alleviate over-smoothing, thus effectively enhancing inference quality. 3) DuetGraph achieves SOTA performance on both inductive and transductive KG reasoning tasks, with up to an 8.7% improvement in quality and a 1.8$\times$ acceleration in training efficiency.
@@ 2 Background
A knowledge graph (KG) is a structured representation of information where entities are represented as nodes, and the relationships between these entities are represented as edges. Typically, a KG $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{R}\}$ is composed of: a set of entities $\mathcal{V}$, a set of relations $\mathcal{R}$, and a set of triplets =$\mathcal{E}=\{(h_{i},r_{i},t_{i})\mid h_{i},t_{i}\in\mathcal{V},r_{i}\in\mathcal{R}\}$, where each triplet represents a directed edge $h_{i}\xrightarrow{r_{i}}t_{i}$ between a head entity $h_{i}$ and a tail entity $t_{i}$.
@@
Given a KG $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, KG completion is to infer and predict missing elements within triplets to enrich the knowledge graph. Depending on the missing component, the task can be categorized into three subtypes: head entity completion (?$,r,t)$, tail entity completion $(h,r,?)$, and relation completion $(h,?,t)$. In this paper, we primarily focus on tail entity completion, following the setting of recent KG works [17, 18], as the other tasks can be reformulated into this one.
@@
KG reasoning methods can be classified based on their use of structural information: message passing-based methods, which primarily leverage local structures, and transformer-based methods, which mainly exploit global structures. Message passing-based methods, such as [3, 25, 26], suffer from well-known limitations of message-passing networks, including incompleteness [27] and over-squashing [28]. Transformer-based methods, such as [29, 30], also have drawbacks. For example, they typically transform graph structures into sequential representations during knowledge encoding, potentially losing critical structural information inherent to KGs [31]. Hybrid approaches that combine message-passing and transformers, such as [17, 18], leverage the strengths of both paradigms. However, they still face key challenges in effectively integrating and balancing local features learned via message passing with global KG information captured by self-attention. Besides, there are also triplet-based methods, such as TransE [32], ComplEx [33], DistMult [34], and RotatE [35], which treat triples as independent instances and often ignore graphs’ topological structure. Beyond these categories, other approaches include meta-learning methods like MetaSD [36], rule path–based models like RNNLogic [37], and tensor decomposition methods such as TuckER-IVR [38]. However, the optimization process of these methods does not directly take into account the issue of score over-smoothing. In response, we propose DuetGraph, explicitly addressing the challenge of score over-smoothing in KG reasoning.
@@ 3 DuetGraph
This section introduces DuetGraph, as shown in Figure $2$In dual-pathway model, after obtaining the representations in Step , we employ an MLP to transform each representation into a score. Loss function is defined for each training triplet $(h,r,t)$: $\mathcal{L}=-log(\sigma(t|h,r))-\sum_{t^{\prime}}log(1-\sigma(t^{\prime}|h,r))$, where $\sigma(\cdot|h,r)$ denotes the score of a candidate triplet, $t^{\prime}$ denotes negative samples. Negative samples are generated by masking the correct answer and uniformly sampling with replacement from the remaining unmasked entities [39]. Finally, we update model parameters by optimizing negative sampling loss [35] with the Adam optimizer [40].. The core architecture of DuetGraph consists of two components: a dual-pathway model for training (Steps ①-④), and a coarse-to-fine reasoning optimization for inference (Steps ⑤-⑧). Section 3.1 presents the dual-pathway global-local fusion model, and Section 3.2 details the coarse-to-fine reasoning optimization.
@@ 3.2 Coarse-to-Fine Reasoning Optimization
In this stage, we first obtain an entity-to-score table by using the coarse model (Step ). Then, the table is split into two subtables based on their rankings (Step ): a high-score subtable made up of the top-k entities, and a low-score subtable containing the remaining ones. The formal description of this process is as follows. Given a query $(h,r,?)$, let $\mathcal{T}=\{(v,s_{v})\mid v\in\mathcal{V},s_{v}\}$ denote the full entity-to-score table, where $s_{v}$ is the score of entity $v$. $\mathrm{Rank}(v)$ denotes the rank of entity $v$ in descending order of the scores. Accordingly, we split $\mathcal{T}$ into two subtables as follows:
@@
At this stage, we firstly update $\mathcal{T}^{\mathrm{high}}$ and $\mathcal{T}^{\mathrm{low}}$ with the fine model (i.e., the dual-pathway model introduced in Section 3). Then, we extract the entities with the highest score from each subtables, denoted as $(e_{h},s_{e_{h}})$ for $\mathcal{T}^{\mathrm{high}}$ and $(e_{l},s_{e_{l}})$ for $\mathcal{T}^{\mathrm{low}}$. After that, we compute the difference $\gamma=s_{e_{l}}-s_{e_{h}}$ based on the pre-defined threshold $\Delta$. If $\gamma$ exceeds $\Delta$, this indicates that the highest-scoring entity in the low-score subtable clearly surpasses the highest-scoring entity in the high-score subtable. Therefore, we select the highest-scoring entity from $\mathcal{T}^{\mathrm{low}}$ as the final answer; otherwise, we select that from $\mathcal{T}^{\mathrm{high}}$.
@@ 4 Empirical Evaluation
In this section, we conduct extensive experiments to answer the following research questions: (RQ1) Can DuetGraph effectively improve the performance of inductive KG reasoning tasks? (RQ2) Can DuetGraph effectively improve the performance of transductive KG reasoning tasks? (RQ3) Can DuetGraph demonstrate strong scalability in KG reasoning tasks by achieving high training efficiency? (RQ4) How is the effectiveness of the components of DuetGraph? (RQ5) Is DuetGraph sensitive to hyperparameter $k$, where $k$ denotes the number of entities in high-score subset?
@@ 4.2 Performance
To answer (RQ2), we evaluate the performance of DuetGraph on four widely utilized transductive KG reasoning datasets. Table 2 demonstrates the impressive performance of DuetGraph compared to baseline models. Specifically, DuetGraph demonstrates substantial performance gains over baseline methods, with improvements of up to 37.2% in MRR, 52.8% in Hits@1, and 24.1% in Hits@10.
@@
CompGCN [12] 0.355 26.4 53.5 0.479 44.3 54.6 0.463 39.2 59.6 0.456 39.2 56.7
NBFNet [26] 0.364 29.4 58.9 0.494 44.4 57.9 0.531 42.4 64.6 0.543 47.7 70.1
RED-GNN [3] 0.374 28.2 58.9 0.519 46.5 60.2 0.521 44.7 64.1 0.556 47.0 70.7
A*Net [11] 0.373 29.4 59.0 0.544 47.1 61.1 0.537 47.7 66.3 0.570 49.1 72.0
AdaProp [49] 0.389 30.2 59.8 0.540 46.8 61.0 0.554 48.5 67.1 0.558 50.2 72.3
@@
DuetGraph (Ours) 0.456 36.1 62.8 0.594 54.2 70.0 0.593 52.2 71.2 0.632 56.1 74.9
4.3 Efficiency
To answer (RQ3), we evaluate Hits@1 throughout training on the FB15k-237 and YAGO3-10, the latter being a large-scale dataset with millions of training triples. We compare DuetGraph with the best models in each category—AdaProp (message passing-based), SAttLE (transformer-based), and KnowFormer (hybrid). As shown in Figure 3, DuetGraph finally achieves SOTA performance on FB15k-237 while reducing training time by nearly 50% compared to the second best method. We also observe from Figure 3 that DuetGraph achieves SOTA performance on YAGO3-10 while requiring less training time compared to other methods. These results show that DuetGraph achieves scalability through high training efficiency. The observed improvement is primarily attributable to this dual-pathway design, which enables parallel training of both local and global pathways.
@@ 4.4 Ablation Study
We additionally evaluate the performance of DuetGraph on four transductive KG reasoning datasets using three different types coarse-grained reasoning models: a triplet-based model (HousE [19]), a message passing-based model (RED-GNN [3]), and a hybrid message passing-transformer model (KnowFormer [17]), which is the SOTA method. As shown in Table 4, DuetGraph, when integrated with any of the three coarse grained models, consistently outperforms its competitors.
@@
DuetGraph 0.456 36.1 62.8 0.594 54.2 70.0 0.593 52.2 71.2 0.632 56.1 74.9
w/o local 0.445 35.1 61.2 0.584 54.1 69.0 0.582 51 70.3 0.617 54.2 73.7
w/o global 0.441 34.9 61.4 0.565 51.7 66.6 0.586 51 69.8 0.614 53.8 74.4
w/o Coarse-to-Fine reasoning 0.429 33.6 61.0 0.571 52.0 67.3 0.555 48.8 67.4 0.612 54.1 73.9
@@
DuetGraph (w/ triplets-based model as coarse model) 0.456 36.1 62.8 0.594 54.2 70.0 0.593 52.2 71.2 0.632 56.1 74.9
@@ 4.5 Parameter Analysis
To answer (RQ5), we conduct experiments by varying the hyperparameter $k$ introduced in Section 3.2 across four different transductive datasets. As shown in Figure 4, DuetGraph is insensitive to the parameter $k$, suggesting stable performance. We further conduct hyperparameter experiments on inductive datasets, as presented in Appendix D.2, and obtain consistent results.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Knowledge graphs (KGs) are structured representations of real-world entities and their relationships, widely applied in domains such as information retrieval [1, 2], logical reasoning [3, 4], recommendation systems [5, 6], materials science [7, 8], and biomedical research [9, 10]. However, existing KGs are often incomplete, missing certain factual information [11, 12], which limits their effectiveness in downstream applications. As a result, inferring and completing missing entity information through KG reasoning is essential.
@@
However, such a single stage stacking approach tends to result in the problem of score over-smoothing, where incorrect answers receive scores similar to correct ones (Figure 1111The $x$-axis denotes normalized score gap ($\frac{|\text{Score}_{\text{correct}}-\text{Score}_{\text{incorrect}}|}{std(\text{Score})}$) and $y$-axis indicates entity count per interval.), making them hard to distinguish. Accordingly, we summarize the problem into two core challenges. Challenge 1: Existing studies [23, 24] have shown that stacking either message-passing or attention layers individually deepens information propagation and aggravates the over-smoothing problem. When message-passing and attention layers are stacked together, these effects accumulate. Challenge 2: The discriminative capacity of single stage models is typically limited [25], as they generate the answer directly based on a one-shot reasoning. This deficiency in discrimination further exacerbates the over-smoothing phenomenon [26].
@@
[…]
@@ 2 Background
A knowledge graph (KG) is a structured representation of information where entities are represented as nodes, and the relationships between these entities are represented as edges. Typically, a KG $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathcal{R}\}$ is composed of: a set of entities $\mathcal{V}$, a set of relations $\mathcal{R}$, and a set of triplets $\mathcal{E}=\{(h_{i},r_{i},t_{i})\mid h_{i},t_{i}\in\mathcal{V},r_{i}\in\mathcal{R}\}$, where each triplet represents a directed edge $h_{i}\xrightarrow{r_{i}}t_{i}$ between a head entity $h_{i}$ and a tail entity $t_{i}$.
@@
Given a KG $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{R})$, KG completion is to infer and predict missing elements within triplets to enrich the knowledge graph. Depending on the missing component, the task can be categorized into three subtypes: head entity completion $(?,r,t)$, tail entity completion $(h,r,?)$, and relation completion $(h,?,t)$. In this paper, we primarily focus on tail entity completion, following the setting of recent KG works [19, 20], as the other tasks can be reformulated into this one (See Appendix D.2).
@@
KG reasoning methods can be classified based on their use of structural information: message passing-based methods, which primarily leverage local structures, and transformer-based methods, which mainly exploit global structures. Message passing-based methods, such as [5, 27, 28], suffer from well-known limitations of message-passing networks, including incompleteness [29] and over-squashing [30]. Transformer-based methods, such as [31, 32], also have drawbacks. For example, they typically transform graph structures into sequential representations during knowledge encoding, potentially losing critical structural information inherent to KGs [33, 34]. Hybrid approaches that combine message-passing and transformers, such as [19, 20], leverage the strengths of both paradigms. However, they still face key challenges in effectively integrating and balancing local features learned via message passing with global KG information captured by self-attention. Besides, there are also triplet-based methods, such as TransE [35], ComplEx [36], DistMult [37], and RotatE [38], which treat triples as independent instances and often ignore graphs’ topological structure. Beyond these categories, other approaches include meta-learning methods like MetaSD [31], rule path–based models like RNNLogic [39], and tensor decomposition methods such as TuckER-IVR [40]. However, the optimization process of these methods does not directly take into account the issue of score over-smoothing. In response, we propose DuetGraph, explicitly addressing the challenge of score over-smoothing in KG reasoning.
@@ 3 DuetGraph
This section introduces DuetGraph, as shown in Figure 2222In dual-pathway model, after obtaining the representations in Step , we employ an MLP to transform each representation into a score. Loss function is defined for each training triplet $(h,r,t)$: $\mathcal{L}=-log(\sigma(t|h,r))-\sum_{t^{\prime}}log(1-\sigma(t^{\prime}|h,r))$, where $\sigma(\cdot|h,r)$ denotes the score of a candidate triplet, $t^{\prime}$ denotes negative samples. Negative samples are generated by masking the correct answer and uniformly sampling with replacement from the remaining unmasked entities [41]. Finally, we update model parameters by optimizing negative sampling loss [38] with the Adam optimizer [42].. The core architecture of DuetGraph consists of two components: a dual-pathway model for training (Steps ①-④), and a coarse-to-fine reasoning optimization for inference (Steps ⑤-⑧). Section 3.1 presents the dual-pathway global-local fusion model, and Section 3.2 details the coarse-to-fine reasoning optimization.
@@ 3.2 Coarse-to-Fine Reasoning Optimization
In this stage, we first obtain an entity-to-score table by using the coarse model (Step ). Then, the table is split into two subtables based on their rankings (Step ): a high-score subtable made up of the top‑k entities, and a low-score subtable containing the remaining ones. The formal description of this process is as follows. Given a query $(h,r,?)$, let $\mathcal{T}=\{(v,s_{v})\mid v\in\mathcal{V},s_{v}\}$ denote the full entity-to-score table, where $s_{v}$ is the score of entity $v$. $\mathrm{Rank}(v)$ denotes the rank of entity $v$ in descending order of the scores. Accordingly, we split $\mathcal{T}$ into two subtables as follows:
@@
At this stage, we firstly update $\mathcal{T}^{\mathrm{high}}$ and $\mathcal{T}^{\mathrm{low}}$ with the fine model (i.e., the dual-pathway model introduced in Section 3). Then, we extract the entities with the highest score from each subtables, denoted as $(e_{h},s_{e_{h}})$ for $\mathcal{T}^{\mathrm{high}}$ and $(e_{l},s_{e_{l}})$ for $\mathcal{T}^{\mathrm{low}}$. After that, we compute the difference $\gamma=s_{e_{l}}-s_{e_{h}}$ based on the pre-defined threshold $\Delta$. If $\gamma$ exceeds $\Delta$, this indicates that the highest-score entity in the low-score subtable clearly surpasses the highest-score entity in the high-score subtable. Therefore, we select the highest-score entity from $\mathcal{T}^{\mathrm{low}}$ as the final answer; otherwise, we select that from $\mathcal{T}^{\mathrm{high}}$.
@@ 4 Empirical Evaluation
In this section, we conduct extensive experiments to answer the following research questions: (RQ1) Can DuetGraph effectively improve the performance of inductive KG reasoning tasks? (RQ2) Can DuetGraph effectively improve the performance of transductive KG reasoning tasks? (RQ3) Can DuetGraph demonstrate strong scalability in KG reasoning tasks by achieving high training efficiency? (RQ4) How is the effectiveness of the components of DuetGraph? (RQ5) In the coarse-to-fine reasoning, what is the standard of the coarse model? (RQ6) Is DuetGraph sensitive to hyperparameter $k$, where $k$ denotes the number of entities in a high-score subset? (RQ7) How generalizable is DuetGraph across tasks on knowledge graphs?
@@ 4.1 Experiments setup
Triple Classification Datasets.
For the triple classification task, we conduct experiments on three widely used knowledge graph datasets: UMLS[52], FB13[53] and WN11[53].
Triple Classification Baselines.
The following four categories of SOTA models are adopted as baselines for comparison with DuetGraph in triple classification: triplet-based (HousE [21]), message passing-based (AdaProp [51]), transformer-based (HittER [54]), and hybrid message passing-transformer (KnowFormer [19]) models (SOTA methods for comprehensive comparison).
@@ 4.2 Performance
To answer (RQ2), we evaluate the performance of DuetGraph on four widely utilized transductive KG reasoning datasets. Table 2 demonstrates the impressive performance of DuetGraph compared to baseline models. Specifically, DuetGraph demonstrates substantial performance gains over baseline methods, with improvements of up to 37.1% in MRR, 52.8% in Hits@1, and 24.0% in Hits@10.
@@
CompGCN [14] 0.355 26.4 53.5 0.479 44.3 54.6 0.463 38.3 59.6 0.421 39.2 57.7
NBFNet [28] 0.415 32.1 59.9 0.551 49.7 66.6 0.525 45.1 63.9 0.563 48.0 70.8
RED-GNN [5] 0.374 28.3 55.8 0.533 48.5 62.4 0.543 47.6 65.1 0.556 48.3 68.9
A*Net [13] 0.411 32.1 58.6 0.549 49.5 65.9 0.521 44.7 63.1 0.556 47.0 70.7
AdaProp [51] 0.417 33.1 58.5 0.562 49.9 67.1 0.554 49.3 65.5 0.573 51.0 68.5
@@
DuetGraph (Ours) 0.453 36.1 62.4 0.593 54.2 69.9 0.590 52.1 71.2 0.631 56.1 74.8
4.3 Efficiency.
To answer (RQ3), we evaluate Hits@1 throughout training on the FB15k-237 and YAGO3-10, the latter being a large-scale dataset with millions of training triples. We compare DuetGraph with the best models in each category—AdaProp (message passing-based), SAttLE (transformer-based), and KnowFormer (hybrid). As shown in Figure 3, DuetGraph finally achieves SOTA performance on FB15k-237 while reducing training time by nearly 50% compared to the second-best method. We also observe from Figure 3 that DuetGraph achieves SOTA performance on YAGO3-10 while requiring less training time compared to other methods. These results show that DuetGraph achieves scalability through high training efficiency. The observed improvement is primarily attributable to this dual-pathway design, which enables parallel training of both local and global pathways.
@@
Furthermore, to demonstrate the efficiency of DuetGraph on very large KGs, we conduct experiments on Wikidata5M [61] and Freebase [62], and compare DuetGraph with highly efficient rule-based methods (e.g., AnyBURL [63]). It is worth noting that for the Wikidata5M and Freebase datasets, the AnyBURL paper [63] does not explicitly specify under which data split the results in Table 2 of [63] were obtained. For a fair comparison, we use the same data split as in [62].
Table 3: Comparison of different methods across very large knowledge graphs. (The best results are bolded in red with a yellow highlight.)
Method Wikidata5M Freebase
MRR H@1 H@10 Learning Inference MRR H@1 H@10 Learning Inference
AnyBURL (Rule-based) 0.350 30.9 42.9 10,000s 4,462s 0.588 53.6 68.2 10,000s 3,142s
KnowFormer (Emb-based) 0.332 26.7 46.3 31,436s 105s 0.684 65.7 73.6 32,109s 176s
DuetGraph (Emb-based) 0.363 32.7 49.5 28,866s 80s 0.697 69.3 73.8 30,158s 141s
As shown in Table 3, DuetGraph achieves SOTA quality performance with a learning time in the same order of magnitude as AnyBURL [63], and it demonstrates a significant reduction in inference time compared to AnyBURL [63]. It demonstrates that DuetGraph maintains high efficiency and strong quality even when applied to extremely large knowledge graphs.
@@ 4.4 Ablation Study
To answer (RQ5), we additionally evaluate the performance of DuetGraph on four transductive KG reasoning datasets using three different types of coarse-grained reasoning models: a triplet-based model (HousE [21]), a message passing-based model (RED-GNN [5]), and a hybrid message passing-transformer model (KnowFormer [19]). The triplet-based model focuses exclusively on local triple-level patterns. The message passing model captures neighborhood-level information, and the transformer model handles global patterns (like our fine model).
As shown in Table 5, DuetGraph, when integrated with any of the three coarse-grained models, consistently outperforms its competitors. Among them, the triplet-based model achieves the best performance as a coarse model. This aligns with prior work [64], which shows that maximizing architectural diversity between coarse and fine models leads to better overall performance; in this case, the triplet-based model benefits from its maximal architectural difference from our global-information-focused fine model.
@@
DuetGraph 0.453 36.1 62.4 0.593 54.2 69.9 0.590 52.1 71.2 0.631 56.1 74.8
w/o local 0.445 35.1 61.2 0.584 54.1 69.0 0.582 51.0 70.3 0.617 54.2 73.7
w/o global 0.441 34.9 61.4 0.565 51.7 66.6 0.586 51.0 69.8 0.614 53.8 74.4
w/o Coarse-to-Fine reasoning 0.437 34.8 61.1 0.580 53.0 68.9 0.567 50.5 67.7 0.616 54.8 73.5
@@
DuetGraph (w/ triplets-based model as coarse model) 0.453 36.1 62.4 0.593 54.2 69.9 0.590 52.1 71.2 0.631 56.1 74.8
@@ 4.5 Parameter Analysis
To answer (RQ6), we conduct experiments by varying the hyperparameter $k$ introduced in Section 3.2 across four different transductive datasets. As shown in Figure 4, DuetGraph is insensitive to the parameter $k$, suggesting stable performance. We further conduct hyperparameter experiments on inductive datasets, as presented in Appendix D.4, and obtain consistent results.
@@
4.6 Generalization.
To answer (RQ7), we evaluate DuetGraph’s performance on the triple classification task. The experimental results in Table 6 demonstrate that DuetGraph consistently outperforms all baseline methods across all datasets, achieving new SOTA performance on the triple classification task, which further highlights the task generality of our proposed DuetGraph framework.
Table 6: Comparison of triple classification accuracy on different datasets. (The best results are bolded in red with a yellow highlight.)
Method UMLS Acc (%) FB13 Acc (%) WN11 Acc (%)
HousE [21] 83.1 69.8 65.3
HittER [54] 59.4 62.2 69.6
AdaProp [51] 77.0 71.9 67.1
KnowFormer [19] 83.1 77.3 70.2
DuetGraph 84.3 80.0 71.9